OpenStudy (anonymous):

what is the derivative of x^2/(x^2-9)

OpenStudy (anonymous):

quotient rule for this one \[\left(\frac{f}{g}\right)'=\frac{gf'-fg'}{g^2}\] with \[f(x)=x^2, f'(x)=2x,g(x)=x^2-9,g'(x)=2x\] plug and chug

OpenStudy (mathmale):

I'd suggest you begin by asking yourself what kind of function you have here; that will dictate therule of differentiation that you'll need to use.

OpenStudy (anonymous):

yeah i know that and this is what i got -18x/(x^2-9)^2 and i,m wondering if its wrong